# Hiding Local Manipulations on SAR Images: a Counter-Forensic Attack

### v1

@@ V-B Manipulated Dataset
V-C Counter-forensic attack parameters
V-C1 Image despeckling
V-C2 SAR System Frequency Response
V-D Detector setup
To evaluate the strength of our counter-forensic attack, we employed two different versions of the forensic detectors proposed in [10], namely the SAR Adapted Extractor (SAE) and the Augmented SAR Adapted Extractor (ASAE), which are slightly different one with the other according to the specific training pipeline employed. Specifically, the SAE detector is trained as described in Section II-C, while the ASAE detector considers data augmentation in training phase, applying upscaling to the input images to enlarge the set of available SAR products (see [10] for more details).
In our experiments, we used the detectors in the best possible conditions, i.e., we trained and tested them over a homogenous dataset, in which the same product was available both in training and testing stages. In doing so we guaranteed that, in absence of counter-forensic attacks, the detectors performed at the best of their possibilities. Indeed, we wanted to avoid the risk of introducing errors due to different data distributions in train and test stages. This methodology ensured that the unique factor which might cause a degradation in the performance is the proposed counter-forensic attack.
To train and test the forensic detectors in the best possible conditions, we always showed in training phase all the possible semantic scenarios: for each of the four considered geographical areas (Rome, Dublin, Gaza and Italy mountains), we randomly selected $70\%$ of data for training (further split in $60-40\%$ for training and validation sets, respectively) and left the remaining $30\%$ for testing. Notice that the training images were $100\%$ pristine, i.e., the detectors have not been trained to spot local image forgeries. Contrarily, as done in [10], the images of the test set did always present local manipulations. This operation allows to test the detection performances in presence of tampering: if this is not the case (i.e., the test image is pristine), the detectors would report an almost flat fingerprint due to absence of inconsistencies [10]. The tampering operations were realized as explained in V-B. For each amplitude image in the test set, we considered one possible splicing area, performing on the donor image all the editing operations listed in Table I.
Following similar considerations, we stopped the detectors’ pipeline at the fingerprint extraction step, i.e., we did not include the mask estimation step. The motivation is that extracting only the fingerprint and not the mask allows to reduce potential errors (with consequent performance drop) introduced during the binarization of the mask. As specified before, we put ourselves in the best possible testing conditions, in order to evaluate the amount of performance degradation cause by the suggested methodology.
@@ VI-A Quality of the attacked images
Image quality metrics. As done in [12], we compare pristine and attacked samples by exploiting several similarity metrics [^2] In the supplemental materials, we also compare attacked and pristine images for what concerns their frequency behaviour.. We evaluate the visual quality of the attacked samples by computing metrics typical of the image full-reference quality assessment field, i.e., the Structural Similarity Index Measure (SSIM) [34], and the Multi-Scale Structural Similarity Index Measure (MS-SSIM) [35]. The higher the similarity measure between attacked and pristine samples, ideally approaching $1$, the better the quality of the attack, in terms of producing pixel content that visually resembles original data.
We also consider SAR compatibility metrics to evaluate our ability in reproducing speckle pattern in the attacked images. In particular, we rely on the Equivalent Number of Looks (ENL) [36], usually employed in the SAR despeckling field, to understand if the considered attacked samples are compatible with the original SAR data. The ENL is a statistical measure to quantify the level of speckle noise in SAR images. In particular, we compute the absolute relative difference of the ENL of attacked images with respect to their original pristine version. We define this metrics as $|\Delta\text{-}\textrm{ENL}|$. Our goal is achieving low values of $|\Delta\text{-}\textrm{ENL}|$, meaning for a realistic reproduction of speckle noise in the attacked samples.
Despeckling influence. We start evaluating the achieved similarity metrics according to presence or not of the despeckling step in our attack. To do so, we restrict our investigations to one of the four situations presented in Section IV-B2, i.e., the different levels of attacker knowledge on the SAR system frequency response $\mathbf{H}$. In particular, we omit the case in which $\mathbf{H}$ is fully known, as we have no access to all the information describing the original SAR system used to capture the considered dataset. We consider the best case scenario, which reasonably consists in having available the original complex image $\mathbf{I}e^{j\mathbf{\Phi}}$ under attack.
TABLE II: Quality metrics (SSIM / MSSIM / $|\Delta\text{-}\textrm{ENL}|$) between pristine amplitude images and their attacked versions in presence or absence of despeckling. We hypothesize the attacker has available the original complex image $\mathbf{I}e^{j\mathbf{\Phi}}$ for estimating the frequency response $\mathbf{H}$. In bold, the best result per column.
@@
SSIM / MSSIM / $|\Delta\text{-}\textrm{ENL}|$ SSIM / MSSIM / $|\Delta\text{-}\textrm{ENL}|$ SSIM / MSSIM / $|\Delta\text{-}\textrm{ENL}|$
$\mathbf{H}_{G}$ $\mathbf{0.985}$ / $\mathbf{0.997}$ / $306.2\%$ $\mathbf{0.983}$ / $\mathbf{0.997}$ / $414.5\%$ $\mathbf{0.991}$ / $\mathbf{0.998}$ / $0.518\%$
$\mathbf{H}_{R}$ $0.984$ / $0.997$ / $\mathbf{305.2}\%$ $0.983$ / $0.997$ / $\mathbf{412.7}\%$ $0.991$ / $0.998$ / $0.471\%$
$\mathbf{H}_{D}$ $0.985$ / $0.997$ / $305.7\%$ $0.983$ / $0.997$ / $413.7\%$ $0.991$ / $0.998$ / $\mathbf{0.458\%}$
@@
TABLE III: Quality metrics (SSIM / MSSIM / $|\Delta\text{-}\textrm{ENL}|$) between pristine amplitude images and their attacked versions. In bold, the best results per column.
@@
SSIM / MSSIM / $|\Delta\text{-}\textrm{ENL}|$ SSIM / MSSIM / $|\Delta\text{-}\textrm{ENL}|$ SSIM / MSSIM / $|\Delta\text{-}\textrm{ENL}|$ SSIM / MSSIM / $|\Delta\text{-}\textrm{ENL}|$
$\mathbf{H}_{G}$ $\mathbf{0.991}$ / $\mathbf{0.998}$ / $0.518\%$ $\mathbf{0.991}$ / $\mathbf{0.998}$ / $0.506\%$ $\mathbf{0.991}$ / $\mathbf{0.998}$ / $0.519\%$
$\mathbf{H}_{R}$ $0.991$ / $0.998$ / $0.471\%$ $0.991$ / $0.998$ / $\mathbf{0.446}\%$ $0.991$ / $0.998$ / $\mathbf{0.449}\%$
$\mathbf{H}_{D}$ $0.991$ / $0.998$ / $\mathbf{0.458\%}$ $0.991$ / $0.998$ / $0.462\%$ $0.991$ / $0.998$ / $0.462\%$ $\mathbf{0.989}$ / $\mathbf{0.998}$ / $\mathbf{8.847}\%$
The achieved results of image processing and SAR data compatibility metrics are depicted in Table II. SSIM and MS-SSIM do not show significant differences between the possible scenarios: in all situations, SSIM and MS-SSIM exceed $0.98$ and $0.997$ respectively, meaning for excellent visual quality. Nonetheless, omitting the despeckling step allows to achieve the best performance.
@@ VI-B Counter-forensic Attack Evaluation
TABLE IV: Average AUC in absence of counter-forensic attack. In bold, the highest achieved AUC per editing operation.
Editing operation
Detector Gaussian Blur Upscale Near Upscale Far Downscale Near Downscale Far Rotate Near Rotate Far
SAE $0.933$ $0.941$ $0.947$ $0.829$ $\mathbf{0.681}$ $0.924$ $0.918$
ASAE $\mathbf{0.974}$ $\mathbf{0.988}$ $\mathbf{0.996}$ $\mathbf{0.841}$ $0.659$ $\mathbf{0.967}$ $\mathbf{0.948}$
TABLE V: AUC in case of counter-forensic attack. First, we show the absolute AUC value, then, the relative difference with respect to the AUC in absence of attack (see Table IV). In bold, the lowest achieved AUC per editing operation and the maximum deviation with respect to the AUC in absence of attack. Better results are associated with lower AUCs, ideally approaching $0.5$, even though AUCs $\approx 0.6/0.7$ already achieve valid performances.
Editing operation
Gaussian Upscale Upscale Downscale Downscale Rotate Rotate
Detector Blur Near Far Near Far Near Far
@@
ASAE $\mathbf{0.621}/\mathbf{-36.24}\%$ $\mathbf{0.636}/\mathbf{-35.63}\%$ $\mathbf{0.674}/\mathbf{-32.33}\%$ $\mathbf{0.586}/\mathbf{-30.32}\%$ $\mathbf{0.557}/-15.48\%$ $\mathbf{0.616}/\mathbf{-36.30}\%$ $\mathbf{0.611}/\mathbf{-35.55}\%$
Tampering localization metrics. For each analyzed image, we evaluate the tampering localization performances by comparing the detector output fingeprint $\mathbf{P}$ and the groundtruth tampering mask $\mathbf{M}$. In particular, we rely on the Area Under the Curve (AUC) of the Receiver Operating Characteristic (ROC) curve of the related binary classification problem. In absence of counter-forensic attacks, we expect the AUC of the ROC to approach $1$, meaninig for high detector accuracy in spotting local manipulations. Our goal is to reduce the AUC, ideally achieving $0.5$, which means the detector is not anymore able to distinguish between pristine and manipulated pixels. Indeed, we want the fingerprints of attacked manipulated images to resemble those of pristine data, i.e., to be overall homogeneous in the entire pixel area [^4] In the supplemental materials, we also compare the detector fingerprints of attacked manipulated and pristine images for what concerns their pixel values distribution, showing that they report similar characteristics..
Detector results in absence of attack. To give an upper bound on the achievable detection performances, we first provide the results on the manipulated images not undergone to any counter-forensic attacks. Table IV depicts the achieved results, averaged over the test dataset, by varying the editing operation applied to the donor image.
@@
Detector results in case of attack. Table V depicts the average detection results in case of testing manipulated amplitude images undergone the proposed counter-forensic attack. It is worth noticing an important performance degradation in all scenarios. The highest achieved AUC is $0.728$, in contrast with the $0.996$ achieved in absence of the attack.
We can observe another interesting fact: the most sensitive detector to the attack is the ASAE, which is actually the best performing detector in standard situations. This implies that, if the forensic analyst is not aware of potential global attacks done to hide the manipulation area, it is quite probable that the selected detector will be the ASAE and that localization results in case of attack will be worse than those achieved by the SAE. Nonetheless, even in case of SAE detector, we report a strong performance degradation with respect to normal conditions. Apart from the “Downscale Far” editing which was already challenging to detect before the attack, all the editing operations undergo AUC degradation of more than $23\%$ and $30\%$ for SAE and ASAE detectors, respectively.
@@
((a)) The AUC before the attack is $0.991$; the AUC after the attack is $0.622$.
((b)) The AUC before the attack is $0.984$; the AUC after the attack is $0.639$.
Figure 9: Examples of the proposed counter-forensic attack applied to manipulated amplitude images. From left to right: the manipulated image in absence of counter-forensic attack; the manipulated image in presence of counter-forensic attack; the ASAE detector fingerprint in absence of counter-forensic attack; the ASAE detector fingerprint in presence of counter-forensic attack.
((a)) The AUC before the attack is $0.967$; the AUC after the attack is $0.545$.
((b)) The AUC before the attack is $0.930$; the AUC after the attack is $0.599$.
Figure 10: Examples of the proposed counter-forensic attack applied to manipulated amplitude images by exploiting military vehicles selected from MSTAR dataset. From left to right: the donor SAR amplitude image of the vehicle; the manipulated image in presence of attack; close-up of the pristine image of the area inside the white box; close-up of the manipulated image in absence of attack; close-up of the manipulated image in presence of attack; close-up of the ASAE detector fingerprint in absence of attack; close-up of the ASAE detector fingerprint in presence of attack.
VI-C Realistic Attacks
In this section, we test the proposed methodology on particular cases which resemble challenging real-world attacks. Indeed, the previous analysis allows a large number of precisely controlled experiments, though squared forgery areas which have been automatically generated might be little realistic.
We consider an additional dataset including SAR data known as Moving and Stationary Target Acquisition and Recognition (MSTAR), which contains public X-band complex SAR images of military tank targets at $1$ foot resolution ($\sim 30\textrm{cm}$) [37]. We create local manipulations by inserting target vehicles in SAR amplitude images selected from our TerraSAR-X testing dataset. Given that the average size of a military tank is at maximum $10\times 4$ meters in length and witdh, we consider donor pixel areas around $35\times 16$ pixels. This size is not fixed as it depends on the specific target vehicle to insert.

### v2

@@ V-B Manipulated Dataset
Figure 7: Examples of the considered SAR amplitude data. First column depicts pristine images. Second column shows manipulated version of the same images; the splicing area is contained inside the red box. Third column shows a close-up of the forgery area and of the surrounding pixels (white box). Last column depicts a close-up with the proposed counter-forensic attack. Since the considered images have $16$-bit dynamics, we show their amplitudes in logarithmic scale.
V-C Counter-forensic attack parameters
V-C1 Image despeckling
V-C2 SAR System Frequency Response
@@
V-D Detectors’ setup
In our experiments, we used the detectors in the best possible conditions, i.e., we trained and tested them over a homogeneous dataset, in which the same product was available both in training and testing stages. In doing so we guaranteed that, in absence of counter-forensic attacks, the detectors performed at the best of their possibilities. Indeed, we wanted to avoid the risk of introducing errors due to different data distributions in train and test stages. This methodology ensured that the unique factor which might cause a degradation in the performance is the proposed counter-forensic attack.
To train and test the forensic detectors in the best possible conditions, we always showed in training phase all the possible semantic scenarios: for each of the four considered geographical areas (Rome, Dublin, Gaza and Italy mountains), we randomly selected $70\%$ of data for training (further split in $60-40\%$ for training and validation sets, respectively) and left the remaining $30\%$ for testing.
[…]
The images of the test set did always present local manipulations. This operation allows to test the detection performances in presence of tampering: if this is not the case (i.e., the test image is pristine), the detectors would report an almost flat heatmap due to absence of inconsistencies [10, 32]. The tampering operations were realized as explained above.
@@ VI-A Quality of the attacked images
Image quality metrics. As done in [12], we compare pristine and attacked samples by exploiting several similarity metrics [^3] In the supplemental materials, we also compare attacked and pristine images for what concerns their frequency behaviour.. We evaluate the visual quality of the attacked samples by computing metrics typical of the image full-reference quality assessment field, i.e., the Peak Signal to Noise Ratio (PSNR) and the Multi-Scale Structural Similarity Index Measure (MS-SSIM) [40]. The higher these metrics, the better the quality of the attack, in terms of producing pixel content that visually resembles original data. Notice that the PSNR has to be intended in the uint16 dynamic range, thus we expect higher values with respect to those typically occurring in counter-forensics for uint8 images [41, 15, 16, 19].
We also consider SAR compatibility metrics to evaluate our ability in reproducing speckle pattern in the attacked images. In particular, we rely on the Equivalent Number of Looks (ENL) [42], usually employed in the SAR despeckling field, to understand if the considered attacked samples are compatible with the original SAR data. The ENL is a statistical measure to quantify the level of speckle noise in SAR images. In particular, we compute the absolute relative difference of the ENL of attacked images with respect to their original pristine version. We define this metrics as $|\Delta\text{-}\textrm{ENL}|=|(\textrm{ENL}_{pristine}-\textrm{ENL}_{attacked})/\textrm{ENL}_{pristine}|$. Our goal is achieving low values of $|\Delta\text{-}\textrm{ENL}|$, meaning for a realistic reproduction of speckle noise in the attacked samples.
Despeckling influence. We start evaluating the achieved metrics according to presence or not of the despeckling step in our attack. To do so, we restrict our investigations to one of the four situations presented in Section IV-B2, i.e., the different levels of attacker knowledge on the SAR system frequency response $\mathbf{H}$. In particular, we omit the case in which $\mathbf{H}$ is fully known, as we have no access to all the information describing the original SAR system used to capture the considered dataset. We consider the best case scenario, which reasonably consists in having available the original complex image $\mathbf{I}e^{j\mathbf{\Phi}}$ under attack.
TABLE II: Quality metrics (PSNR [dB] / MSSIM / $|\Delta\text{-}\textrm{ENL}|$) between pristine amplitude images and their attacked versions in presence or absence of despeckling. We hypothesize the attacker has available the original complex image $\mathbf{I}e^{j\mathbf{\Phi}}$ for estimating the frequency response $\mathbf{H}$. In bold, the best result per column.
@@
PSNR / MSSIM / $|\Delta\text{-}\textrm{ENL}|$ PSNR / MSSIM / $|\Delta\text{-}\textrm{ENL}|$ PSNR / MSSIM / $|\Delta\text{-}\textrm{ENL}|$
$\mathbf{H}_{G}$ $\mathbf{50.90}$ / $\mathbf{0.997}$ / $306.2\%$ $\mathbf{50.65}$ / $\mathbf{0.997}$ / $414.5\%$ $\mathbf{52.64}$ / $\mathbf{0.998}$ / $0.518\%$
$\mathbf{H}_{R}$ $50.83$ / $0.997$ / $\mathbf{305.2}\%$ $50.59$ / $0.997$ / $\mathbf{412.7}\%$ $52.46$ / $0.998$ / $0.471\%$
$\mathbf{H}_{D}$ $50.87$ / $0.997$ / $305.7\%$ $50.62$ / $0.997$ / $413.7\%$ $52.55$ / $0.998$ / $\mathbf{0.458\%}$
@@
TABLE III: Quality metrics (PSNR [dB] / MSSIM / $|\Delta\text{-}\textrm{ENL}|$) between pristine amplitude images and their attacked versions. In bold, the best results per column.
@@
PSNR / MSSIM / $|\Delta\text{-}\textrm{ENL}|$ PSNR / MSSIM / $|\Delta\text{-}\textrm{ENL}|$ PSNR / MSSIM / $|\Delta\text{-}\textrm{ENL}|$ PSNR / MSSIM / $|\Delta\text{-}\textrm{ENL}|$
$\mathbf{H}_{G}$ $\mathbf{52.64}$ / $\mathbf{0.998}$ / $0.518\%$ $\mathbf{52.67}$ / $\mathbf{0.998}$ / $0.506\%$ $\mathbf{52.64}$ / $\mathbf{0.998}$ / $0.519\%$
$\mathbf{H}_{R}$ $52.46$ / $0.998$ / $0.471\%$ $52.43$ / $0.998$ / $\mathbf{0.446}\%$ $52.47$ / $0.998$ / $\mathbf{0.449}\%$
$\mathbf{H}_{D}$ $52.55$ / $0.998$ / $\mathbf{0.458\%}$ $52.53$ / $0.998$ / $0.462\%$ $52.56$ / $0.998$ / $0.462\%$ $\mathbf{52.39}$ / $\mathbf{0.998}$ / $\mathbf{8.847}\%$
The achieved results of image processing and SAR data compatibility metrics are depicted in Table II. PSNR and MS-SSIM do not show significant differences between the possible scenarios: in all situations, PSNR and MS-SSIM exceed $50$ dB and $0.997$ respectively, meaning for excellent visual quality. Nonetheless, omitting the despeckling step allows to achieve the best performance.
@@ VI-B Counter-forensic Attack Evaluation
TABLE IV: Average AUC, BA, F1 and IoU in absence of counter-forensic attack. In bold, the highest metric achieved per editing operation.
DownScaleFar DownScaleNear GaussianBlur-0.5 RotateFar RotateNear UpScaleFar UpScaleNear
AUC / BA / F1 / IoU AUC / BA / F1 / IoU AUC / BA / F1 / IoU AUC / BA / F1 / IoU AUC / BA / F1 / IoU AUC / BA / F1 / IoU AUC / BA / F1 / IoU
SAE $0.98/0.97/0.92/0.89$ $1.0/0.99/0.97/0.94$ $1.0/1.0/0.99/0.97$ $1.0/1.0/0.98/0.97$ $1.0/1.0/0.98/0.97$ $1.0/1.0/0.99/0.98$ $1.0/1.0/0.99/0.98$
ASAE $0.99/0.98/0.94/0.91$ $0.99/0.99/0.97/0.95$ $\mathbf{1.0/1.0/0.99/0.98}$ $\mathbf{1.0/1.0/0.99/0.98}$ $\mathbf{1.0/1.0/0.99/0.98}$ $\mathbf{1.0/1.0/1.0/0.99}$ $\mathbf{1.0/1.0/0.99/0.99}$
TruForS $\mathbf{1.0/1.0/0.96/0.93}$ $\mathbf{1.0/1.0/0.97/0.95}$ $1.0/1.0/0.98/0.97$ $1.0/1.0/0.98/0.96$ $1.0/1.0/0.98/0.96$ $1.0/1.0/0.99/0.98$ $1.0/1.0/0.99/0.97$
Mean $0.99/0.98/0.94/0.91$ $1.0/0.99/0.97/0.95$ $1.0/1.0/0.99/0.97$ $1.0/1.0/0.98/0.97$ $1.0/1.0/0.98/0.97$ $1.0/1.0/0.99/0.98$ $1.0/1.0/0.99/0.98$
TABLE V: Average AUC, BA, F1 and IoU in case of counter-forensic attack. In bold, the lowest metric achieved per editing operation.
DownScaleFar DownScaleNear GaussianBlur-0.5 RotateFar RotateNear UpScaleFar UpScaleNear
AUC / BA / F1 / IoU AUC / BA / F1 / IoU AUC / BA / F1 / IoU AUC / BA / F1 / IoU AUC / BA / F1 / IoU AUC / BA / F1 / IoU AUC / BA / F1 / IoU
SAE $0.8/0.78/0.46/0.40$ $0.76/0.74/0.37/0.32$ $0.79/0.77/0.44/0.39$ $0.79/0.77/0.44/0.38$ $0.78/0.76/0.42/0.37$ $0.88/0.87/0.67/0.61$ $0.81/0.79/0.49/0.43$
ASAE $\mathbf{0.69/0.68}/0.29/0.24$ $\mathbf{0.68/0.68}/0.27/0.24$ $\mathbf{0.72/0.71}/0.35/0.31$ $\mathbf{0.71/0.7}/0.33/0.29$ $\mathbf{0.71/0.71}/0.34/0.3$ $\mathbf{0.81/0.80}/0.54/0.49$ $\mathbf{0.75/0.74}/0.41/0.37$
TruForS $0.77/0.72/\mathbf{0.22/0.15}$ $0.74/0.70/\mathbf{0.20/0.13}$ $0.78/0.73/\mathbf{0.24/0.16}$ $0.78/0.73/\mathbf{0.24/0.16}$ $0.76/0.72/\mathbf{0.23/0.15}$ $0.89/0.84/\mathbf{0.47/0.35}$ $0.81/0.76/\mathbf{0.29/0.20}$
Mean $0.75/0.73/0.32/0.26$ $0.73/0.71/0.28/0.23$ $0.76/0.74/0.34/0.29$ $0.76/0.73/0.34/0.28$ $0.75/0.73/0.33/0.27$ $0.86/0.84/0.56/0.48$ $0.79/0.76/0.40/0.33$
TABLE VI: Average relative difference of all the metrics (over all editing operations), between the case of counter-forensic attack and the absence of it. In bold, the worst case among the three detectors.
AUC BA F1 IoU
SAE $-19.62\%$ $-21.26\%$ $-51.78\%$ $-56.77\%$
ASAE $\mathbf{-27.37\%}$ $\mathbf{-27.99\%}$ $-63.27\%$ $-67.09\%$
TruForS $-21.0\%$ $-25.71\%$ $\mathbf{-72.46\%}$ $\mathbf{-80.72\%}$
Mean $-22.66\%$ $-24.99\%$ $-62.50\%$ $-68.19\%$
Tampering localization metrics. For each analyzed image, we evaluate the tampering localization performances by comparing the detectors output heatmap ($\bar{\mathbf{M}}$ or $\tilde{\mathbf{M}}$ in absence or presence of attack, respectively) and the groundtruth tampering mask $\mathbf{M}$. In particular, we rely on several metrics: the Area Under the Curve (AUC), the Balanced Accuracy (BA), the F1 score (F1) and the Intersection Over Union (IoU). For the metrics requiring a thresholding operation on the heatmap (i.e., the BA, the F1 and the IoU), we always select their maximum value. We do this to evaluate the detectors in their best possible conditions. In absence of counter-forensic attacks, we expect all metrics to approach $1$, meaning for high detectors accuracy in spotting local manipulations. Our goal is to reduce them, ideally achieving AUC and BA around $0.5$ and IoU and F1 around 00, which means the detectors are not anymore able to distinguish between pristine and manipulated pixels.
Detectors results in absence of attack. To give an upper bound on the achievable detection performances, we first provide the results on the manipulated images not undergone to any counter-forensic attacks. Table IV depicts the achieved results, averaged over the test dataset, by varying the editing operation applied to the donor image. The metrics associated with all editing operations always achieve high scores, especially for the ASAE detector, which is known to be more accurate than the SAE [10], and for the TruForS detector.
Detector results in case of attack. Table V depicts the average detection results in case of testing manipulated amplitude images undergone the proposed counter-forensic attack. It is worth noticing an important performance degradation in all scenarios. To clarify, we also report in Table VI the relative metrics difference with respect to the absence of attack, averaged across all the editing operations. On average, the AUC and the BA drop by $\sim 25\%$, while the F1 and the IoU by more than $60\%$. We can observe another interesting fact: the most sensitive detectors to the attack are the ASAE and the TruForS, which are actually the best performing detectors in standard situations. This implies that, if the forensic analyst is not aware of potential global attacks done to hide the manipulation area, it is quite probable that the selected detector will be either the ASAE or the TruForS and that localization results in case of attack will be worse than those achieved by the SAE. Nonetheless, even in case of SAE detector, we report a strong performance degradation with respect to normal conditions.
@@
((a)) Before: AUC/BA/F1$=0.97/0.96/0.89$. After: AUC/BA/F1$=0.65/0.62/0.09$.
((b)) Before: AUC/BA/F1$=0.94/0.94/0.90$. After: AUC/BA/F1$=0.67/0.63/0.08$.
((c)) Before: AUC/BA/F1$=0.94/0.95/0.91$. After: AUC/BA/F1$=0.72/0.64/0.24$.
Figure 8: Examples of the proposed counter-forensic attack applied to manipulated amplitude images. From left to right: the manipulated image in absence of counter-forensic attack; the manipulated image in presence of counter-forensic attack; the TruForS detector heatmap in absence of counter-forensic attack; the TruForS detector heatmap in presence of counter-forensic attack.
VI-C Comparison with state of the art
In this section, we perform further experiments to validate the need of the proposed speckle injection step composing our counter-forensic pipeline. Specifically, we compare our attack with another counter-forensic solution that has been successfully employed to fool tampering localization detectors on natural images [19]. This method is based on the hypothesis that high-frequency patterns carry the most significant forensic traces. Consequently, it focuses on removing high-frequency traces, effectively acting as a denoising step to deceive forensic detectors.
We select this methodology based on the following considerations: (i) it does not require a dedicated training phase, unlike [15, 16]; (ii) it does not rely on prior knowledge of the targeted detector’s functioning, in contrast to [15, 16, 17, 18]; (iii) it does not require access to the distribution of pristine samples, unlike [16, 18]; (iv) it does not explicitly necessitate an adaptation phase for SAR data. These characteristics enable a direct comparison with our proposed attack without requiring modifications. Furthermore, this method has demonstrated highly promising counter-forensic effects against the TruFor detector on natural images, reducing its localization performance by more than $24\%$ across various datasets [19].
To evaluate its effectiveness, we first verify the quality of attacked images, then, we evaluate its counter-forensic effect.
@@ VI-C Comparison with state of the art
Notably, the method proposed in [19] provides better PSNR and MS-SSIM than our proposed attack, even though the results are comparable. Nonetheless, as it was already pointed out, a more discriminant metrics for SAR-related data is the $|\Delta\text{-}\textrm{ENL}|$, being the ENL a statistical measure to quantify the level of speckle noise in images. Our proposed approach returns a $|\Delta\text{-}\textrm{ENL}|=0.458\%$, enabling to better reduce the difference between the ENL of attacked images and that of pristine ones. We believe that method [19], being focused on removing high-frequency traces and actually acting as a denoising step, partially fails in reproducing the unique granular properties of the speckle noise.
Counter-forensic attack evaluation. Table VII compares the average localization results after the counter-forensic attack developed in [19] and our proposed attack. As we can quickly observe, the attack proposed in [19] does not effectively hide manipulation traces in SAR images. Indeed, the difference between localization metrics before and after the attack is negligible. In contrast, our proposed methodology effectively conceals these traces, leading to a significant performance drop across all metrics.
TABLE VII: Localization results, averaged over all detectors (SAE, ASAE and TruForS) and over all editing operations, achieved on images attacked with the method proposed in [19] and with our proposed method, in terms of average AUC, BA, F1 and IoU. We show results after the attack and the relative difference with respect to the absence of attack ($\Delta\%$). In bold, the best counter-forensic results.
AUC BA F1 IoU
attack / $\Delta\%$ attack / $\Delta\%$ attack / $\Delta\%$ attack / $\Delta\%$
@@
Ours $\mathbf{0.77/-22.66\%}$ $\mathbf{0.75/-24.99\%}$ $\mathbf{0.37/-62.50\%}$ $\mathbf{0.31/-68.19\%}$
VI-D Realistic Attacks
In this section, we test the proposed methodology on particular cases which resemble even more challenging real-world attacks. Indeed, the previous analysis allows a large number of precisely controlled experiments, though squared forgery areas which have been automatically generated might be little realistic.
We consider an additional dataset including SAR data known as Moving and Stationary Target Acquisition and Recognition (MSTAR), which contains public X-band complex SAR images of military tank targets at $1$ foot resolution ($\sim 30\textrm{cm}$) [43]. We create local manipulations by inserting target vehicles in SAR amplitude images selected from our TerraSAR-X testing dataset. Given that the average size of a military tank is at maximum $10\times 4$ meters in length and width, we consider donor pixel areas around $35\times 16$ pixels. This size is not fixed as it depends on the specific target vehicle to insert.